# Rapidly-rotating Population III stellar models as a source of primary nitrogen

### v1

@@ Abstract
Context. The first stars might have been fast rotators. This would have important consequences for their radiative, mechanical and chemical feedback.
@@
Methods. We explore the evolution of Population III stars with initial masses in the range of 9 $M_{\odot}$ $\leqslant M_{\rm ini}\leqslant$ 120 $M_{\odot}$ starting with an initial rotation on the Zero Age Main Sequence equal to 70% of the critical one.
Results. We find that with the physics of rotation considered here, our rapidly-rotating Population III stellar models do not follow a homogeneous evolution. They lose very little mass in case mechanical winds are switched on when the surface rotation becomes equal or larger than the critical velocity. Impact on the ionising flux appears modest when compared to moderately-rotating models. Fast rotation favours, in models with initial masses above $\sim$20 $M_{\odot}$, the appearance of a very extended intermediate convective zone around the H-burning shell during the core He-burning phase. This shell has important consequences on the sizes of the He- and CO-cores and thus impacts the final fate of stars. Moreover, it has a strong impact on nucleosynthesis boosting the production of primary 14N.
Conclusions. Fast initial rotation impacts significantly the chemical feedback of Population III stars. Observations of extremely metal-poor stars and/or starbursting regions are essential to provide constraints on the properties of the first stars.
@@ 1 Introduction
According to cosmological simulations, the first stars formed at redshifts $z\sim 20-35$ (Abel et al., 2002; Bromm et al., 2002; Yoshida et al., 2003; Ohkubo et al., 2009; Hirano et al., 2014; Susa, 2019) and were responsible for the first stellar radiative, chemical, and mechanical feedback in the Universe (see e.g. the recent review by Klessen & Glover, 2023, and references therein). To know their evolution is therefore important for questions such as reionisation (Bromm et al., 2002; Choudhury & Ferrara, 2005; Yoon et al., 2012; Sibony et al., 2022), the first chemical enrichments (Meynet & Maeder, 2002b; Prantzos, 2003; Chiappini et al., 2005a; Meynet et al., 2006; Tominaga et al., 2007; Chiappini et al., 2008; Rollinde et al., 2009; Greif et al., 2010; Pallottini et al., 2014; Marassi et al., 2015; Sarmento et al., 2017; Hirai et al., 2018, 2019; Corazza et al., 2022; Sanati et al., 2023), the origin of Carbon-Enhanced Metal-Poor (CEMP) stars (Ji et al., 2015; Clarkson et al., 2018; Liu et al., 2021b; Zepeda et al., 2023), the Population (Pop) III progenitors of long Soft Gamma Ray Bursts (Wang et al., 2012) and of magnetorotational supernovae (Suwa et al., 2007; Mészáros & Rees, 2010; Suwa & Ioka, 2011; de Souza et al., 2011; Maeder & Meynet, 2012; Piro et al., 2014) or of intermediate mass and/or merging black holes (Inayoshi et al., 2016, 2017; Regan et al., 2020; Kinugawa et al., 2020, 2021; Liu et al., 2021a; Tanikawa et al., 2021; Santoliquido et al., 2023; Iorio et al., 2023), and for setting the stage for the next star formation events. In addition, recent JWST observations of high-redshift galaxies (e.g. GLASS and CEERS surveys) have discovered a subset of N-rich galaxies (see e.g. Isobe et al., 2023). In particular the intriguing $\sim$11 source GN-z11 (Cameron et al., 2023), is suggested to contain a sub-population of Pop III stars (Maiolino et al., 2023). Such Pop III stars might be very interesting candidates for enriching in a short timescales these high redshift galaxies in nitrogen.
@@
The stellar feedback processes depend on characteristics of the first stellar populations, such as the initial mass function (Bromm et al., 1999; Abel et al., 2002; Bromm et al., 2002), the initial distribution of rotations that can contain a large fraction of very rapidly-rotating stars (Stacy et al., 2011, 2013; Hirano & Bromm, 2018a) if magnetic braking is inefficient in the absence of strong magnetic fields (Hirano & Bromm, 2018b; Hirano & Machida, 2022; Kimura et al., 2023), and the degree of multiplicity (see e.g. Stacy et al., 2010; Stacy & Bromm, 2013; Liu et al., 2021a) to cite only a few important features. Simulations of Pop III star formation suggest that Pop III stars had large masses, higher than 100 $M_{\odot}$ due to the lack of cooling mechanism that prevents further fragmentation Bromm et al. (2002). However, the exact mass range is still debated. Recent studies allow the formation of Pop III stars with masses down to 1 $M_{\odot}$ and even lower (Yoshida et al., 2006; Clark et al., 2011; Greif et al., 2011; Susa, 2013; Stacy et al., 2016; Wollenberg et al., 2020; Sugimura et al., 2020; Park et al., 2021; Prole et al., 2022; Latif et al., 2022).
@@
Models are also needed to find observable signatures of these primordial stars when seen as a population in high redshift star-forming regions (Schaerer, 2002, 2003; Raiter et al., 2010; Ma et al., 2017; Trussler et al., 2023) or when they explode in energetic, very luminous transient events (Joggerst & Whalen, 2011; Toma et al., 2011; Mesler et al., 2014; Tolstov et al., 2016; Lee et al., 2023). Pop III stars are also privileged objects for studying some frontiers of physics. Since they are born in dark matter mini halos they may be impacted by the physics of the dark matter (Freese et al., 2008; Taoso et al., 2008; Natarajan et al., 2009; Schleicher et al., 2009; Spolyar et al., 2009; Choplin et al., 2017; Liu et al., 2019a, b, 2020). Also, some cosmological theories predict that fundamental constants, such as the fine structure constant, may evolve as a function of time along the cosmic history. Pop III stellar models have been used to test the impact of different values of the fine structure constant on the stellar nucleosynthesis in the very early Universe (Ekström et al., 2010; Huang et al., 2019; Mori & Nomoto, 2020).
@@
In the last two decades or so, evolutionary features of non-rotating Pop III stellar models have been studied among others by Marigo et al. (2001); Chieffi et al. (2001, 2002b, 2002a); Ohkubo et al. (2009); Bahena & Klapp (2010); Heger & Woosley (2010). Works have also been done studying the impact of close binary evolution of Pop III stars (see e.g. Chen et al., 2015; Tsai et al., 2023). The impact of convective interactions between H- and He-burning regions in massive Pop III stellar models has been studied by Clarkson & Herwig (2021).
@@
Ekström et al. (2008) explored the impact of rotation on the evolution and the chemical feedback of Pop III stars. Murphy et al. (2021a) performed a similar study as Ekström et al. (2008), but adopting a different setting for the diffusion coefficients describing the mixing of the elements by rotation. More recently, Aryan et al. (2023) using MESA computed the evolution of Pop III rotating 25 $M_{\odot}$ models.
@@
Rotation deeply impacts nucleosynthesis. For instance, Meynet & Maeder (2002b); Chiappini et al. (2006); Ekström et al. (2008) have shown that the diffusion of He-burning products into the H-burning shell induced by shear turbulence boosts the production of isotopes such as 13C and 14N. Further works have shown that the diffusion of nitrogen into the He-burning core boosts the production of 19F, 22Ne, s- and p-process elements (Meynet & Maeder, 2002a; Meynet et al., 2006; Chiappini et al., 2008; Pignatari et al., 2008; Cescutti et al., 2013; Frischknecht et al., 2016; Limongi & Chieffi, 2018; Choplin et al., 2018, 2022) and give them a status (at least in part) of primary elements. Let us recall that a primary element is an element whose production does not scale with the initial metallicity but is produced from the transformation of hydrogen and helium (see e.g. the book by Matteucci, 2001). An interesting example is the production of primary nitrogen in metal-poor rotating stars (Meynet & Maeder, 2002a). The turbulence induced by rotation triggers mixing between the H- and He-burning zones, allowing some carbon and oxygen produced in the He-burning zone to diffuse into the H-burning zone. In the H-burning shell, carbon and oxygen are transformed into nitrogen by the action of the CNO cycle. This nitrogen is qualified as primary because it is produced from carbon and oxygen produced by the star and not from carbon and oxygen that were present in the star at its formation.
@@
In past works, it has been shown that for a given initial mass, metallicity, and rotation, the primary nitrogen production only occurs at low metallicity, typically at metallicities between $Z\sim 10^{-8}$ and $10^{-4}$, where $Z$ is the mass fraction of the abundances of all the elements heavier than helium (see e.g. Meynet et al., 2006). The reason why this mixing is efficient only at low metallicities is mainly due to the fact that the lack of CNO element at the beginning of the evolution makes the H-burning occur at higher temperatures than in more metal-rich stars. As a consequence, the H- and He-burning zones are less distant from each other in metal-poor stars, and the entropy gradients are weaker, making any mixing either by convection or by diffusion easier.
@@
Previous modelling of zero metallicity stars has suggested that Pop III stars will produce significantly less primary nitrogen than a model with a very small initial metal content (assuming models starting with a velocity at the Zero Age Main Sequence (ZAMS), $\upsilon_{\rm ini}$, corresponding to the same constant fraction of the critical velocity [^1] The critical velocity is the velocity at the equator where the gravity is balanced by the centrifugal force., $\upsilon_{\rm crit}$, at all metallicities). This occurs because some degree of differential rotation is required to trigger the mixing. The differential rotation is mainly produced by the contraction of the core at the end of the Main-Sequence (MS) phase. Contraction in a Pop III star is not as great as in an even slightly metal-enriched star, because as explained before, the temperature of He-burning is not much greater than the temperature in the H-burning core. A modest core contraction is therefore sufficient for raising the core temperatures at the level needed for achieving radiative equilibrium. Furthermore, the opacity in the envelope is lower facilitating the evacuation of the energy released by contraction from the envelope. Thus, less energy is used to expand the envelope, which also suppresses differential rotation. The star remains in the blue part of the Hertzsprung-Russel diagram (HRD) for the whole or at least a substantial part of the core He-burning phase [^2] Farrell et al. (2022) has performed very interesting numerical experiments showing how artificially preventing the contraction actually prevents the crossing of the Hertzsprung-Russel (HR) gap. Actually, things are more complex in the sense that some contraction always occurs and other factors such as the distribution of helium around the H-burning shell are important too..
@@
In this paper, we explore Pop III stellar models rotating faster on the ZAMS than in our previous works and how significant the production of primary nitrogen can be in those. Therefore, instead of considering an initial velocity of $\upsilon_{\rm ini}=0.4\,\upsilon_{\rm crit}$ as in Murphy et al. (2021a), we shall consider $\upsilon_{\rm ini}=0.7\,\upsilon_{\rm crit}$. In Sect. 2, we describe the ingredients of the stellar models. Some effects of fast rotation on the evolution of primordial stars are discussed in Sect. 3. A brief presentation of the small grid computed in this work is given in Sect. 4. We discuss the impact of fast rotation on the primary nitrogen production in Sect. 5. Sect. 6 compares the yields obtained from models at different metallicities, obtained by different authors. The impact of the present yields in a simple chemical evolution model is also presented. The conclusions are given in Sect. 7.
@@ 2 Ingredients of stellar models
The rapidly-rotating Pop III models have been computed with the Geneva stellar evolution code (Genec, see a detailed description in Eggenberger et al., 2008) with the same physical ingredients as the Pop III models published by Murphy et al. (2021a) and models at higher metallicities by Ekström et al. (2012); Georgy et al. (2013a); Groh et al. (2019); Eggenberger et al. (2021); Yusof et al. (2022). The reader is invited to refer to those papers for the details of the input physics. Only a few relevant ingredients are recalled here.
@@
The initial composition considered is $X=0.7516$, $Y=0.2484$, and $Z=0$, with $X$, $Y$, and $Z$ the mass fractions of hydrogen, helium, and heavy elements, respectively. The primordial 4He abundance is coming from predicted values from the Big Bang nucleosynthesis (Cyburt et al., 2003).
@@
We explore the effects of fast rotation on the evolution of massive stars with similar masses as considered in Murphy et al. (2021a). We restricted here to masses between 9 and 120 $M_{\odot}$. More massive star models have been recently explored using Genec in Martinet et al. (2023) and Nandal et al. (2023a, c) although not with fast rotation.
@@
Rotation is implemented according to the stellar-rotation theory developed by Zahn (1992). The expressions describing the vertical and horizontal shear diffusion coefficients are respectively from Maeder (1997) and Chaboyer & Zahn (1992). Magnetic fields are not accounted for in the present models.
@@
We did not account for any mass loss due to line-driven winds, as it was discussed in Murphy et al. (2021a). Even though, there are many other processes that they may allow Pop III stars to lose mass (see e.g. the discussion in the introduction of Liu et al., 2021b), for example, pulsation-driven mass loss (Volpato et al., 2023), mass loss during a mass transfer in a close binary and mechanical mass loss, in this work only mechanical mass loss has been taken into account. Mechanical mass loss occurs when the surface rotation reaches the critical velocity. We estimate the mass lost mechanically as the mass that needs to be removed from the star in order to keep its surface rotation subcritical. A detailed account for the exact numerical procedure is explained in Georgy et al. (2013b, Sect. 2.2).
@@
The critical velocity is defined as the surface rotation such that the centrifugal acceleration at the equator balances the gravity there. In this study, we followed the expression given by Maeder & Meynet (2000) which accounts for the fact that the critical velocity varies when the luminosity of the star is near the Eddington luminosity (the so-called $\Omega\Gamma$ limit where $\Omega$ is the angular velocity and $\Gamma$ the Eddington factor). The critical rotation has two roots depending on the Eddington factor:
@@
$\upsilon_{\rm crit}=\begin{cases}\sqrt{\frac{2}{3}\frac{GM}{R_{\text{p,crit}}}}&,\text{$\Gamma<0.693$}\\ \sqrt{\frac{3}{2}\upsilon_{\rm crit,1}\frac{R_{\text{e,crit}}}{R_{\text{p,crit}}}\frac{1-\Gamma}{{V}^{\prime}(\omega))}}&,\text{$\Gamma>0.693$}\\ \end{cases}$ (1)
@@
where $R_{\text{p,crit}}=\frac{2}{3}R_{\text{e,crit}}$, $R_{\text{p,crit}}$ and $R_{\text{e,crit}}$ are the polar and equatorial radius at critical velocity respectively, $V^{\prime}(\omega)$ is the ratio between the actual volume of the deformed rotating star to a spherical volume with a radius equal to $R_{\text{p,crit}}$, and $\omega$ is the ratio of the surface angular velocity to the critical angular velocity (given by the first root in the equation just above).
@@
The limits of the convective cores have been determined using the Schwarzschild criterion, with a step overshooting parameter of $d_{\rm over}/H_{P}=0.1$ for the core H- and He-burning phase ($d_{\rm over}$ is the spatial extension of the convective core above the radius given by the Schwarzschild’s limit and $H_{P}$ is the pressure scale height at the Schwarzschild’s boundary).
@@
We complement the grid by Murphy et al. (2021a) for specific masses (9, 20, 60, 85, and 120 $M_{\odot}$) with models having an initial rotation rate of $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.7$. This gives initial rotations on the ZAMS that are between 650 and 1200 km/s (see column 2 in Table 4333Keeping constant values for $\upsilon_{\rm ini}/\upsilon_{\rm crit}$ for all the initial masses implies a faster initial surface velocity for more massive stars.). At first sight, these rotations may appear quite extreme values, at least well outside the range of surface velocities measured in the present-day Universe. However, the perception is different if we assume that stars begin their evolution on the ZAMS with a total angular momentum content that is kept constant for all metallicities. Since Pop III stars of a given mass have a much smaller momentum of inertia than solar metallicity stars, they naturally have faster rotations. To give a numerical example, the total angular momentum of a solar metallicity 9 $M_{\odot}$ model with an initial surface velocity equal to 40% the critical one ($264\,{\rm km}/{\rm s}$) is $0.072\times 10^{53}\,{\rm g}\,{\rm cm}^{2}/{\rm s}$ (models by Ekström et al., 2012). The total angular momentum of a Pop III 9 $M_{\odot}$ model with an initial surface velocity equal to 70% the critical one ($649\,{\rm km}/{\rm s}$) is $0.085\times 10^{53}\,{\rm g}\,{\rm cm}^{2}/{\rm s}$. The angular momenta are not very different, but the rotation rates are.

### v2

# Rapidly rotating Population III stellar models as a source of primary nitrogen

@@ Abstract
Context. The first stars might have been fast rotators. This would have important consequences for their radiative, mechanical, and chemical feedback.
@@
Methods. We explore the evolution of Population III stars with initial masses in the range of 9 $M_{\odot}$ $\leqslant M_{\rm ini}\leqslant$ 120 $M_{\odot}$, starting with an initial rotation on the zero-age main sequence equal to 70% of the critical one.
Results. We find that with the physics of rotation considered here, our rapidly rotating Population III stellar models do not follow a homogeneous evolution. They lose very little mass in the case in which mechanical winds are switched on when the surface rotation becomes equal to or larger than the critical velocity. The impact on the ionising flux appears to be modest when compared to moderately rotating models. Fast rotation favours, in models with initial masses above $\sim$20 $M_{\odot}$, the appearance of a very extended intermediate convective zone around the H-burning shell during the core He-burning phase. This shell has important consequences for the sizes of the He- and CO-cores, and thus impacts the final fate of stars. Moreover, it has a strong impact on nucleosynthesis, boosting the production of primary 14N.
Conclusions. Fast initial rotation significantly impacts the chemical feedback of Population III stars. Observations of extremely metal-poor stars and/or starbursting regions are essential to provide constraints on the properties of the first stars.
@@ 1 Introduction
According to cosmological simulations, the first stars formed at redshifts of $z\sim 20-35$ (Abel et al., 2002; Bromm et al., 2002; Yoshida et al., 2003; Ohkubo et al., 2009; Hirano et al., 2014; Susa, 2019) and were responsible for the first stellar radiative, chemical, and mechanical feedback in the Universe (see e.g. the recent review by Klessen & Glover, 2023, and references therein). Knowing how they evolved is therefore important for questions such as reionisation (Bromm et al., 2002; Choudhury & Ferrara, 2005; Yoon et al., 2012; Sibony et al., 2022), the first chemical enrichments (Meynet & Maeder, 2002b; Prantzos, 2003; Chiappini et al., 2005a; Meynet et al., 2006; Tominaga et al., 2007; Chiappini et al., 2008; Rollinde et al., 2009; Greif et al., 2010; Pallottini et al., 2014; Marassi et al., 2015; Sarmento et al., 2017; Hirai et al., 2018, 2019; Corazza et al., 2022; Sanati et al., 2023), the origin of carbon-enhanced metal-poor (CEMP) stars (Ji et al., 2015; Clarkson et al., 2018; Liu et al., 2021b; Zepeda et al., 2023), the Population (Pop) III progenitors of long soft gamma-ray bursts (Wang et al., 2012), magnetorotational supernovae (SNe) (Suwa et al., 2007; Mészáros & Rees, 2010; Suwa & Ioka, 2011; de Souza et al., 2011; Maeder & Meynet, 2012; Piro et al., 2014), or intermediate-mass and/or merging black holes (Inayoshi et al., 2016, 2017; Regan et al., 2020; Kinugawa et al., 2020, 2021; Liu et al., 2021a; Tanikawa et al., 2021; Santoliquido et al., 2023; Iorio et al., 2023), and for setting the stage for the next star formation events. In addition, recent JWST observations of high-redshift galaxies (e.g. the GLASS and CEERS surveys) have discovered a subset of N-rich galaxies (see e.g. Isobe et al., 2023). In particular, it has been suggested that the intriguing $\sim$11 source GN-z11 (Cameron et al., 2023) contains a sub-population of Pop III stars (Maiolino et al., 2023). Such Pop III stars might be very interesting candidates for enriching these high-redshift galaxies with nitrogen over short timescales.
@@
The stellar feedback processes depend on characteristics of the first stellar populations, such as the initial mass function (Bromm et al., 1999; Abel et al., 2002; Bromm et al., 2002), the initial distribution of rotations, which can contain a large fraction of very rapidly rotating stars (Stacy et al., 2011, 2013; Hirano & Bromm, 2018a) if magnetic braking is inefficient in the absence of strong magnetic fields (Hirano & Bromm, 2018b; Hirano & Machida, 2022; Kimura et al., 2023), and the degree of multiplicity (see e.g. Stacy et al., 2010; Stacy & Bromm, 2013; Liu et al., 2021a), to cite only a few important features. Simulations of Pop III star formation suggest that Pop III stars had large masses higher than 100 $M_{\odot}$ due to the lack of a cooling mechanism that would prevent further fragmentation Bromm et al. (2002). However, the exact mass range is still debated. Recent studies allow the formation of Pop III stars with masses down to 1 $M_{\odot}$ and even lower (Yoshida et al., 2006; Clark et al., 2011; Greif et al., 2011; Susa, 2013; Stacy et al., 2016; Wollenberg et al., 2020; Sugimura et al., 2020; Park et al., 2021; Prole et al., 2022; Latif et al., 2022).
@@
Models are also needed to find observable signatures of these primordial stars when seen as a population in high-redshift star-forming regions (Schaerer, 2002, 2003; Raiter et al., 2010; Ma et al., 2017; Trussler et al., 2023) or when they explode in energetic, very luminous transient events (Joggerst & Whalen, 2011; Toma et al., 2011; Mesler et al., 2014; Tolstov et al., 2016; Lee et al., 2024). Pop III stars are also privileged objects for studying some frontiers of physics. Since they are born in dark matter mini halos they may be impacted by the physics of the dark matter (Freese et al., 2008; Taoso et al., 2008; Natarajan et al., 2009; Schleicher et al., 2009; Spolyar et al., 2009; Choplin et al., 2017; Liu et al., 2019a, b, 2020). Also, some cosmological theories predict that fundamental constants, such as the fine structure constant, may evolve as a function of time over cosmic history. Pop III stellar models have been used to test the impact of different values of the fine structure constant on the stellar nucleosynthesis in the very early Universe (Ekström et al., 2010; Huang et al., 2019; Mori & Nomoto, 2020).
@@
In the last two decades or so, evolutionary features of non-rotating Pop III stellar models have been studied by, among others, Marigo et al. (2001); Chieffi et al. (2001, 2002b, 2002a); Ohkubo et al. (2009); Bahena & Klapp (2010); Heger & Woosley (2010). Work has also been done to study the impact of close binary evolution of Pop III stars (see e.g. Chen et al., 2015; Tsai et al., 2023). The impact of convective interactions between H- and He-burning regions in massive Pop III stellar models has been studied by Clarkson & Herwig (2021).
@@
Ekström et al. (2008) explored the impact of rotation on the evolution and chemical feedback of Pop III stars. Murphy et al. (2021a) performed a similar study as Ekström et al. (2008), but adopted a different setting for the diffusion coefficients describing the mixing of the elements by rotation. More recently, Aryan et al. (2023) used MESA to compute the evolution of Pop III rotating 25 $M_{\odot}$ models.
@@
Rotation deeply impacts nucleosynthesis. For instance, Meynet & Maeder (2002b); Chiappini et al. (2006); Ekström et al. (2008) have shown that the diffusion of He-burning products into the H-burning shell induced by shear turbulence boosts the production of isotopes such as 13C and 14N. Further works have shown that the diffusion of nitrogen into the He-burning core boosts the production of 19F, 22Ne, s-, and p-process elements (Meynet & Maeder, 2002a; Meynet et al., 2006; Chiappini et al., 2008; Pignatari et al., 2008; Cescutti et al., 2013; Frischknecht et al., 2016; Limongi & Chieffi, 2018; Choplin et al., 2018, 2022) and given them the status (at least in part) of primary elements. Let us recall that a primary element is an element whose production does not scale with the initial metallicity but is produced from the transformation of hydrogen and helium (see e.g. the book by Matteucci, 2001). An interesting example is the production of primary nitrogen in metal-poor rotating stars (Meynet & Maeder, 2002a). The turbulence induced by rotation triggers mixing between the H- and He-burning zones, allowing some carbon and oxygen produced in the He-burning zone to diffuse into the H-burning zone. In the H-burning shell, carbon and oxygen are transformed into nitrogen by the action of the carbon–nitrogen–oxygen (CNO) cycle. This nitrogen qualifies as primary because it is produced from carbon and oxygen produced by the star and not from carbon and oxygen that were present in the star at its formation.
@@
In past works, it has been shown that for a given initial mass, metallicity, and rotation, the primary nitrogen production only occurs at low metallicity, typically at metallicities between $Z\sim 10^{-8}$ and $10^{-4}$, where $Z$ is the mass fraction of the abundances of all the elements heavier than helium (see e.g. Meynet et al., 2006). The reason why this mixing is efficient only at low metallicities is mainly due to the fact that the lack of CNO elements at the beginning of the evolution makes the H-burning occur at higher temperatures than in more metal-rich stars. As a consequence, the H- and He-burning zones are less distant from each other in metal-poor stars, and the entropy gradients are weaker, making any mixing either by convection or by diffusion easier.
@@
Previous modelling of zero-metallicity stars has suggested that Pop III stars will produce significantly less primary nitrogen than a model with a very small initial metal content (assuming that models start with a velocity at the zero-age main sequence (ZAMS), $\upsilon_{\rm ini}$, corresponding to the same constant fraction of the critical velocity [^1] The critical velocity is the velocity at the equator where the gravity is balanced by the centrifugal force., $\upsilon_{\rm crit}$, at all metallicities). This occurs because some degree of differential rotation is required to trigger the mixing. The differential rotation is mainly produced by the contraction of the core at the end of the main sequence (MS) phase. Contraction in a Pop III star is not as great as in an even slightly metal-enriched star, because as we explained before, the temperature of He-burning is not much greater than the temperature in the H-burning core. A modest core contraction is therefore sufficient to raise the core temperatures to the level needed to achieve radiative equilibrium. Furthermore, the opacity in the envelope is lower, facilitating the evacuation of the energy released by contraction from the envelope. Thus, less energy is used to expand the envelope, which also suppresses differential rotation. The star remains in the blue part of the Hertzsprung-Russel diagram (HRD) for the whole core He-burning phase, or at least a substantial part of it [^2] Farrell et al. (2022) has performed very interesting numerical experiments showing how artificially preventing the contraction actually prevents the crossing of the Hertzsprung-Russel (HR) gap. Actually, things are more complex in the sense that some contraction always occurs and other factors such as the distribution of helium around the H-burning shell are important too..
@@
In this paper, we explore Pop III stellar models rotating faster on the ZAMS than in our previous works and how significant the production of primary nitrogen can be in those. Therefore, instead of considering an initial velocity of $\upsilon_{\rm ini}=0.4\,\upsilon_{\rm crit}$, as in Murphy et al. (2021a), we consider $\upsilon_{\rm ini}=0.7\,\upsilon_{\rm crit}$. In Sect. 2, we describe the ingredients of the stellar models. Some effects of fast rotation on the evolution of primordial stars are discussed in Sect. 3. A brief presentation of the small grid computed in this work is given in Sect. 4. We discuss the impact of fast rotation on the primary nitrogen production in Sect. 5. Sect. 6 compares the yields obtained from models at different metallicities by different authors. The impact of the present yields on a simple chemical evolution model is also presented. The conclusions are given in Sect. 7.
@@ 2 Ingredients of stellar models
The rapidly rotating Pop III models have been computed with the Geneva stellar evolution code (Genec, see a detailed description in Eggenberger et al., 2008) with the same physical ingredients as the Pop III models published by Murphy et al. (2021a) and models at higher metallicities by Ekström et al. (2012); Georgy et al. (2013a); Groh et al. (2019); Eggenberger et al. (2021); Yusof et al. (2022). The reader is invited to refer to those papers for the details of the input physics. Only a few relevant ingredients are recalled here.
@@
The initial composition considered is $X=0.7516$, $Y=0.2484$, and $Z=0$, with $X$, $Y$, and $Z$ the mass fractions of hydrogen, helium, and heavy elements, respectively. The primordial 4He abundance comes from predicted values from the Big Bang nucleosynthesis (Cyburt et al., 2003).
@@
We explore the effects of fast rotation on the evolution of massive stars with similar masses to those considered in Murphy et al. (2021a). We restricted masses here to between 9 and 120 $M_{\odot}$. More massive star models have been recently explored using Genec in Martinet et al. (2023) and Nandal et al. (2023a, b), although not with fast rotation.
@@
Rotation was implemented according to the stellar rotation theory developed by Zahn (1992). The expressions describing the vertical and horizontal shear diffusion coefficients are from Maeder (1997) and Chaboyer & Zahn (1992), respectively. Magnetic fields are not accounted for in the present models.
@@
We did not account for any mass loss due to line-driven winds, as this was discussed in Murphy et al. (2021a). Even though there are many other processes that may allow Pop III stars to lose mass (see e.g. the discussion in the introduction of Liu et al., 2021b) – for example, pulsation-driven mass loss (Volpato et al., 2023), mass loss during a mass transfer in a close binary, and mechanical mass loss – in this work only mechanical mass loss has been taken into account. Mechanical mass loss occurs when the surface rotation reaches a critical velocity. We estimated the mass lost mechanically as the mass that needs to be removed from the star in order to keep its surface rotation sub-critical. A detailed account of the exact numerical procedure is given in Georgy et al. (2013b, Sect. 2.2).
@@
The critical velocity is defined as the surface rotation such that the centrifugal acceleration at the equator balances the gravity there. In this study, we followed the expression given by Maeder & Meynet (2000), which accounts for the fact that the critical velocity varies when the luminosity of the star is near the Eddington luminosity (the so-called $\Omega\Gamma$ limit, where $\Omega$ is the angular velocity and $\Gamma$ the Eddington factor). The critical rotation has two roots depending on the Eddington factor:
@@
$\upsilon_{\rm crit}=\begin{cases}\sqrt{\frac{2}{3}\frac{GM}{R_{\text{p,crit}}}}&,\text{$\Gamma<0.693$}\\ \sqrt{\frac{3}{2}\upsilon_{\rm crit,1}\frac{R_{\text{e,crit}}}{R_{\text{p,crit}}}\frac{1-\Gamma}{{V}^{\prime}(\omega))}}&,\text{$\Gamma>0.693$}\\ \end{cases},$ (1)
@@
where $R_{\text{p,crit}}=\frac{2}{3}R_{\text{e,crit}}$, $R_{\text{p,crit}}$ and $R_{\text{e,crit}}$ are the polar and equatorial radius at critical velocity, respectively, $V^{\prime}(\omega)$ is the ratio between the actual volume of the deformed rotating star to a spherical volume with a radius equal to $R_{\text{p,crit}}$, and $\omega$ is the ratio of the surface angular velocity to the critical angular velocity (given by the first root in the equation just above).
@@
The limits of the convective cores were determined using the Schwarzschild criterion, with a step overshooting parameter of $d_{\rm over}/H_{P}=0.1$ for the core H- and He-burning phase ($d_{\rm over}$ is the spatial extension of the convective core above the radius given by the Schwarzschild’s limit and $H_{P}$ is the pressure scale height at Schwarzschild’s boundary).
@@
We complement the grid by Murphy et al. (2021a) for specific masses (9, 20, 60, 85, and 120 $M_{\odot}$) with models that have an initial rotation rate of $\upsilon_{\rm ini}/\upsilon_{\rm crit}=0.7$. This gives initial rotations on the ZAMS that are between 650 and 1200 km/s (see column 2 in Table 4333Keeping constant values for $\upsilon_{\rm ini}/\upsilon_{\rm crit}$ for all the initial masses implies a faster initial surface velocity for more massive stars.). At first sight, these rotations may appear to have quite extreme values, at least well outside the range of surface velocities measured in the present-day Universe. However, the perception is different if we assume that stars begin their evolution on the ZAMS with a total angular momentum content that is kept constant for all metallicities. Since Pop III stars of a given mass have a much smaller momentum of inertia than solar metallicity stars, they naturally have faster rotations. To give a numerical example, the total angular momentum of a solar metallicity 9 $M_{\odot}$ model with an initial surface velocity equal to 40% of the critical one ($264\,{\rm km}/{\rm s}$) is $0.072\times 10^{53}\,{\rm g}\,{\rm cm}^{2}/{\rm s}$ (models by Ekström et al., 2012). The total angular momentum of a Pop III 9 $M_{\odot}$ model with an initial surface velocity equal to 70% of the critical one ($649\,{\rm km}/{\rm s}$) is $0.085\times 10^{53}\,{\rm g}\,{\rm cm}^{2}/{\rm s}$. The angular momenta are not very different, but the rotation rates are.